1. The specific heat of aluminum is 0.902 J/g•°C. How much heat is absorbed by an aluminum pie tin with a mass of 473 grams to raise its temperature from room temperature (23°C) to oven temperature (190°C)?
  2. Stainless steel accessories in cars are usually plated with chromium to give them a shiny surface and to prevent rusting. When 5 grams of chromium at 23°C absorbs 62.5 J of heat, the temperature increases to 50.8°c. What is the specific heat of chromium?
  3. If 335 g of water at 65.5°C loses 9,750 J of heat, what is the final temperature of the water?
1
Expert's answer
2021-06-25T05:41:49-0400
  1. Q = C x m x deltaT = 0.902 x 473 x (190-23) = 58 451 J
  2. C = Q / (m x deltaT) = 62.5 / (5 x (50.8-23)) = 0.45 J/g•°C
  3. deltaT = Q / (C x m) = 9750 / (4.2 x 335) = 6.9 °C

T2 = 65.5-6.9 = 58.6°C